OpenStudy (anonymous):

Leo brought a bulldozer for 63103. The value of the bulldozer depreciated at a constant rate per year. The table below shows the value of the bulldozer after the first and second year. Year 1 58054.74 Year 2 53410.38 Which function bets represents the value of the bulldozer after t years?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

F(t)=58054.76(0.92)^t F(t)=63103(0.08)^t F(t)= 63103(0.92)^t F(t)=58054.76)(0.08)^t

hero (hero):

You are given a function of the form \(y = ab^x\) where a = the initial value of the bulldozer, b = the constant rate of decrease (x,y) = (t, F(t)) t is in years starting with t = 0. F(t) = Value of the bulldozer after t years.

OpenStudy (anonymous):

Yes

hero (hero):

If you investigate the situation a bit further, you'll notice that you were given depreciation values for the first three years: F(0) = 63103 F(1) = 58054.74 F(2) = 53410.38

hero (hero):

Do you see this?

OpenStudy (anonymous):

Yes

hero (hero):

This means all you have to do is plug in the the values t = {0, 1, 2} in to the correct formula and you should receive the appropriate values for F(t).

hero (hero):

One huge clue is this: If F(0) = 63103, then 63103 must be the initial value of the function. Right?

hero (hero):

I mean if that's the case, we at least know that \(F(t) = (63103)b^t\)

OpenStudy (anonymous):

It's c?

hero (hero):

C is correct.